# Decomposition for Enhancing Attention: Improving LLM-based Text-to-SQL through Workflow Paradigm

### v1

@@ Abstract
In-context learning of large-language models (LLMs) has achieved remarkable success in the field of natural language processing, while extensive case studies reveal that the single-step chain-of-thought prompting approach faces challenges such as attention diffusion and inadequate performance in complex tasks like text-to-SQL. To improve the contextual learning capabilities of LLMs in text-to-SQL, a workflow paradigm method is proposed, aiming to enhance the attention and problem-solving scope of LLMs through decomposition. Specifically, the information determination module for eliminating redundant information and the brand-new prompt structure based on problem classification greatly enhance the model’s attention. Additionally, the inclusion of self-correcting and active learning modules greatly expands the problem-solving scope of LLMs, hence improving the upper limit of LLM-based approaches. Extensive experiments conducted on three datasets demonstrate that our approach outperforms other methods by a significant margin. About 2-3 percentage point improvements compared to the existing baseline on the Spider Dev and Spider-Realistic datasets and new SOTA results on the Spider Test dataset are achieved. Our code is available on GitHub: https://github.com/FlyingFeather/DEA-SQL.
@@
[…]
@@ 1 Introduction
In this paper, we propose a workflow paradigm method for decomposing and enhancing attention based on few-shot. This method draws on human thinking patterns, adheres to the principle of making subtasks as simple as possible, and reduces irrelevant information in each step to specifically enhance the solvable scope of LLM and improve the attention of LLM to enhance their performance. It consists of five sub-modules: an Information Determination module that focuses on attention by reducing interference information through a two-stage method; a Classification & Hint module that solves different problems that cannot be generalized by simply providing different simple prompts; a few-shot SQL Generation module based on question template retrieval; a Self-Correcting module based on error summarization; and an Active Learning module that expands the model’s capabilities based on error cases. To reduce consumption, we minimized the number of few-shots required for each step, with the second stage of field filtering and the self-correction step being zero-shot and the SQL generation module outperforming existing few-shot-based baselines in a zero-shot environment. It has been found that the workflow prompting paradigm is more effective in improving overall performance compared to single-step prompting. In the experiments, execution accuracies of 85.4 and 81.5 were achieved on the Spider-dev and Spider-Realistic datasets, respectively, surpassing existing models and in-context learning schemes.
@@ 2.4 Self-Correction
After analyzing the results generated by the LLM in the previous steps, it is found that although the decomposition in the previous steps reduces redundant information and focuses the effective attention of the LLM, resulting in an improvement in accuracy(The outcome closely aligns with the most recent baseline method, achieving 83.4%. This can be observed in the ’w/0 active & correct’ result in Table 3.) , there are still some issues in field selection, table joining, aggregation, etc. The analysis of 98 incorrect questions reveals the proportion of error types, as depicted in Figure 3. Notably, several significant error points emerge: 1) Extra fields: The LLM often selects an excessive number of fields, rather than limiting its selection to those pertinent to the question; 2) Incorrect fields: For instance, when faced with fields bearing identical names across different tables, the alias may be omitted, leading to errors; 3) Table and field association errors: there may be inconsistencies between the tables and fields used; 4) Fabricated conditions for table joins; 5) Misuse of association words: For example, there is a tendency to habitually use ’left join’ in place of ’join’; 6) Group or order by errors: Mistakes such as incorrect aggregation fields and conditions may be encountered.
@@ 2.5 Active Learning
The model itself has a limited capability threshold and the circles it specializes in do not encompass all problem types in SQL as shown in the Active Learning module in Figure 1. In the experiment, we discovered that the model is more prone to errors for certain problem types (e.g., extremum problems), often producing fixed-paradigm yet incorrect answers. We aim to determine whether the model’s capability thresholds can be expanded through the Active Learning Module to better meet the requirements. The model’s generalization ability is used to present the wrong case and the standard answer to the model to learn the error correction paradigm.
@@
3 Experiment
In this section, we evaluate the proposed algorithm for the text-to-SQL task. Extensive experiments have been conducted to answer the following research questions: RQ1. How does DEA-SQL perform vs. state-of-the-art baselines? RQ2. Whether each module of our method works effectively and how are they impacting the task? RQ3. How do model parameters like the number of few-shots or the information filter layers affect the model?
3.1 Experiment Setup
3.1.1 Datasets
Spider Yu et al. (2018).
Spider Realistic Deng et al. (2021).
3.1.2 Evaluation
3.1.3 Implementation Details
We use OpenAI ChatGPT-3.5 and GPT-4111gpt-3.5-turbo-0613 for ChatGPT and gpt4-0613 for GPT-4. as our base models, which can be assess through official API. For all GPT-based baselines, we utilize the same API to reproduce the models, ensuring fairness in comparison. Our code is available at https://github.com/FlyingFeather/DEA-SQL.
@@ 3.1.3 Implementation Details
3.2 Perfermance (RQ1)
To validate the effectiveness of DEA-SQL in text-to-SQL tasks, a comprehensive comparison was performed on the Spider and Spider-Realistic datasets. We performed most of our experiments on the dev dataset of the Spider dataset, and the experiment was completed just coinciding with the opening of the test set. The experimental results are summarized in Table 1 and 2, where the best performance is shown in boldface. Based on the experimental results, we have made several findings.
@@ 3.3 Ablation Study (RQ2)
To explore the effectiveness of few-shot retrieval methods, we conducted experimental comparisons between different retrieval schemes and zero-shot. In the experiments, it was found that different few-shots have a significant impact on the final accuracy of the results, and choosing poor examples may lead to adverse effects. We explored four retrieval schemes: random, question similarity(ques_sim), template similarity(tem_sim) and template similarity without classification (tem_sim_wo) to find more suitable samples. As shown in Figure 5, random is a little detrimental to overall performance, while the retrieval strategy based on question template similarity in the combined question classification retrieval library yields the best results. Question templates also essentially provide a simple classification for the questions, and relying on question classification methods makes it easier to find the most relevant questions and their solutions, thereby stimulating the capabilities of LLMs.
3.4 Parameters Analysis
In the parameters experiment, we investigated the number of few-shots for the SQL generation step, and the results in Table LABEL:fig:sql_fewshot_nums found that the performance is better when #fewshot=3. This indicates that within a certain range, as the number of effective samples increases, the accuracy of the model improves. The accuracy of the information filter also greatly affects the final accuracy rate. The current method based on LLM has not yet been able to achieve 100% accuracy. It can be seen that with the improvement of the accuracy of the information filter, the final accuracy rate also has a corresponding improvement as shown in Figure LABEL:fig:sql_fc.
@@ 5 Conclusion
Prompting techniques have unleashed the potential of LLMs, with most past research focusing on improving single-step prompting. In the NLP field, structured languages like SQL pose a greater challenge compared to plain text. We introduce the workflow prompting paradigm to enhance the performance of LLMs in text-to-SQL tasks. By decomposing steps, reducing irrelevant information, and transforming the complex problem into a classification task, we can focus the attention of LLMs. Self-correction and active learning modules are designed according to the capability thresholds of LLMs, with different sub-modules performing their respective roles, further improving the upper limit of LLM-based prompting methods in text-to-SQL tasks. Adequate experiments on Spider and Spider-Realistic datasets demonstrate that our approach significantly improves prompting performance, outperforming state-of-the-art fine-tuning and LLM-based methods. This illustrates the effectiveness of the workflow prompting paradigm in enhancing the problem-solving scope of LLMs.
6 Limitations
@@ Appendix B Prompts
B.1 Identify Problem Elements Prompt
@@
B.2 Information filter Prompt
@@ 3.2. The where statement condition only gives the conditions of the corresponding table;
B.3 Question Classification Prompt
@@
B.4 SQL Generation Prompt
B.4.1 few-shot: easy class
@@ B.4 SQL Generation Prompt
B.4.2 few-shot: join class
@@
B.4.3 few-shot: join-nested or nested class
@@
B.5 Self Correction Prompt
@@
B.6 Active Learning Prompt

### v2

@@ Abstract
In-context learning of large-language models (LLMs) has achieved remarkable success in the field of natural language processing, while extensive case studies reveal that the single-step chain-of-thought prompting approach faces challenges such as attention diffusion and inadequate performance in complex tasks like text-to-SQL. To improve the contextual learning capabilities of LLMs in text-to-SQL, a workflow paradigm method is proposed, aiming to enhance the attention and problem-solving scope of LLMs through decomposition. Specifically, the information determination module for eliminating redundant information and the brand-new prompt structure based on problem classification greatly enhance the model’s attention. Additionally, the inclusion of self-correction and active learning modules greatly expands the problem-solving scope of LLMs, hence improving the upper limit of LLM-based approaches. Extensive experiments conducted on three datasets demonstrate that our approach outperforms other methods by a significant margin. About 2-3 percentage point improvements compared to the existing baseline on the Spider Dev, Spider-Realistic, and Bird Dev datasets and new SOTA results on the Spider Test dataset are achieved. Our code is available on GitHub: https://github.com/FlyingFeather/DEA-SQL.
@@
Yuanzhen Xie2   Xinzhou Jin1  Tao Xie2   Mingxiong Lin2  Liang Chen1 Chenyun Yu1  Lei Cheng2  Chengxiang Zhuo2  Bo Hu2  Zang Li2 1Sun Yat-sen University 2Platform and Content Group, Tencent {xieyzh3, taoxie168}@gmail.com jinxzh5@mail2.sysu.edu.cn {chenliang6, yuchy35}@mail.sysu.edu.cn {matrixmxlin, raycheng, felixzhuo, harryyfhu, gavinzli}@tencent.com
@@ 1 Introduction
In this paper, we propose a workflow paradigm method for decomposing and enhancing attention based on few-shot. This method draws on human thinking patterns, adheres to the principle of making subtasks as simple as possible, and reduces irrelevant information in each step to specifically enhance the solvable scope of LLM and improve the attention of LLM to enhance their performance. It consists of five sub-modules: an Information Determination module that focuses on attention by reducing interference information through a two-stage method; a Classification & Hint module that solves different problems that cannot be generalized by simply providing different simple prompts; a few-shot SQL Generation module based on question template retrieval; a Self-Correction module based on error summarization; and an Active Learning module that expands the model’s capabilities based on error cases. To reduce consumption, we minimize the number of few-shots required for each step, with the second stage of field filtering and the self-correction step being zero-shot and the SQL generation module outperforming existing few-shot-based baselines in a zero-shot environment. It has been found that the workflow prompting paradigm is more effective in improving overall performance compared to single-step prompting. In the experiments, execution accuracies of 85.4 and 81.5 were achieved on the Spider-dev and Spider-Realistic datasets, respectively, surpassing existing models and in-context learning schemes.
@@ 2.4 Self-Correction
After analyzing the results generated by the LLM in the previous steps, it is found that although the decomposition in the previous steps reduces redundant information and focuses the effective attention of the LLM, resulting in an improvement in accuracy(The outcome closely aligns with the most recent baseline method, achieving 83.4%. This can be observed in the ’w/0 active & correct’ result in Table 4.), there are still some issues in field selection, table joining, aggregation, etc. The analysis of 98 incorrect questions reveals the proportion of error types, as depicted in Figure 3. Notably, several significant error points emerge: 1) Extra fields: The LLM often selects an excessive number of fields, rather than limiting its selection to those pertinent to the question; 2) Incorrect fields: For instance, when faced with fields bearing identical names across different tables, the alias may be omitted, leading to errors; 3) Table and field association errors: there may be inconsistencies between the tables and fields used; 4) Misuse of association words: For example, there is a tendency to habitually use ’left join’ in place of ’join’; 5) Fabricated conditions for table joins; 6) Group or order by errors: Mistakes such as incorrect aggregation fields and conditions may be encountered.
@@ 2.5 Active Learning
The model itself has a limited capability threshold and the circles it specializes in do not encompass all problem types in SQL as shown in the Active Learning module in Figure 1. In the experiment, we discover that the model is more prone to errors for certain problem types (e.g., extremum problems), often producing fixed-paradigm yet incorrect answers. We aim to determine whether the model’s capability thresholds can be expanded through the Active Learning Module to better meet the requirements. The model’s generalization ability is used to present the wrong case and the standard answer to the model to learn the error correction paradigm.
@@
3 Experiment
In this section, we evaluate the proposed algorithm for the text-to-SQL task. Extensive experiments have been conducted to answer the following research questions: RQ1. How does DEA-SQL perform vs. state-of-the-art baselines? RQ2. Whether each module of our method works effectively and how are they impacting the task? RQ3. What is the token and time cost of the method? RQ4. How do model parameters like the number of few-shots or the information filter layers affect the method?
3.1 Experiment Setup
3.1.1 Implementation Details
We use OpenAI ChatGPT-3.5 and GPT-4111gpt-3.5-turbo-0613 for ChatGPT and gpt4-0613 for GPT-4. as our base models, which can be assessed through official API. For all GPT-based baselines, we utilize the same API to reproduce the models, ensuring fairness in comparison. Our code is available at https://github.com/FlyingFeather/DEA-SQL.
3.1.2 Datasets
Spider Yu et al. (2018).
Spider Realistic Deng et al. (2021).
Bird Li et al. (2024).
Bird is a comprehensive dataset, containing 12,751 distinct question-SQL pairings, spread across 95 large-scale databases that collectively amount to 33.4 GB in size. This dataset covers an expansive range of over 37 professional fields such as blockchain, hockey, healthcare, and education. In a bid to help models generate precise SQL queries, Bird integrates external knowledge as an auxiliary resource. This external knowledge is specifically derived from four sources: numerical reasoning, domain-specific knowledge, synonyms, and value illustration.
3.1.3 Evaluation
3.2 Perfermance (RQ1)
To validate the effectiveness of DEA-SQL in text-to-SQL tasks, a comprehensive comparison was performed on the Spider, Spider-Realistic, and Bird datasets. We performed most of our experiments on the dev dataset of the Spider dataset, and the experiment was completed just coinciding with the opening of the test set. The experimental results are summarized in Table 1, 2 and 3222Due to cost constraints, we only provide the results of DEA-SQL+ChatGPT+zeroshot-SC(reduce Self-Correction module) on Bird Dev dataset, where the best performance is shown in boldface. Based on the experimental results, we have several findings.
@@ 3.2 Perfermance (RQ1)
On the experimental results of the more complex Bird dataset in Table 3, our method still achieved good results, demonstrating considerable scalability. Compared with directly using ChatGPT, we have increased by 15.13 percentage points; compared with MAC-SQL+ChatGPT, we have increased by 1.79 percentage points; even compared with DIN-SQL+GPT4, we have increased by 1.63 percentage points. We firmly believe that applying the method directly to GPT4 will yield greater improvements. Because of the multi-step method, when the accuracy of each step improves, the final accuracy improvement is cumulative.
Table 3: Comparison of previous methods and our method in terms of execution accuracy on the Bird Dev dataset.
Method EX
ChatGPT 37.22
GPT4 46.15
DIN-SQL + GPT-4 50.72
MAC-SQL + GPT-4 50.56
DEA-SQL + ChatGPT+ zeroshot-SC
(reduce Self-Correction module)
52.35
@@ 3.3 Ablation Study (RQ2)
To explore the effectiveness of few-shot retrieval methods, we conducted experimental comparisons between different retrieval schemes and zero-shot. In the experiments, it was found that different few-shots have a significant impact on the final accuracy of the results, and choosing poor examples may lead to adverse effects. We explored four retrieval schemes: random, question similarity(ques_sim), template similarity(tem_sim), and template similarity without classification (tem_sim_wo) to find more suitable samples. As shown in Table 7, random is a little detrimental to overall performance, while the retrieval strategy based on question template similarity in the combined question classification retrieval library yields the best results. Question templates also essentially provide a simple classification for the questions, and relying on question classification methods makes it easier to find the most relevant questions and their solutions, thereby stimulating the capabilities of LLMs.
3.4 Cost Analysis (RQ3)
Table 6: The token and time cost analysis. Inference Time is the average time of running the same case three times using ChatGPT. DEA-SQL-C represents the removal of the classification module in the DEA-SQLmethod.
Method Avg. Token Num Inference Time EX
C3  2803 19.34s 81.2
DIN-SQL  9126 4.37s 83.5
DAIL  700 - 83.6
MAC-SQL  4043 5.45s 78.6
DEA-SQL  5611 4.55s 85.4
DEA-SQL-C  3324 4.16s 83.8
Indeed, the cost is a factor that should be considered in practical use, and we conduct the experiment for an analysis of token and time costs to substantiate the usability of the framework we propose. It can be clearly seen from the Table 6: 1) Our method consumes less in terms of inference time, ensuring the efficiency of our method in real applications. 2) In terms of token consumption, we need some examples for the model to learn from the context, which leads to some increase in tokens. However, as can be seen, it is still more economical than DIN-SQL and brings the best model effect. We believe this is an acceptable trade-off (in addition, the cost and effect balance can be controlled by reducing steps or using traditional solutions to replace some steps). Our core point is the attention-focused workflow paradigm in in-context learning. In practical use scenarios, we can fully decide to replace some modules with other solutions to reduce overall consumption.
3.5 Parameters Analysis (RQ4)
In the parameters experiment, we investigated the number of few-shots for the SQL generation step, and the results in Figure LABEL:fig:sql_fewshot_nums found that the performance is better when #fewshot=3. This indicates that within a certain range, as the number of effective samples increases, the accuracy of the model improves. The accuracy of the information filter also greatly affects the final accuracy rate. The current method based on LLM has not yet been able to achieve 100% accuracy. It can be seen that with the improvement of the accuracy of the information filter, the final accuracy rate also has a corresponding improvement as shown in Figure LABEL:fig:sql_fc.
@@ 5 Conclusion
Prompting techniques have unleashed the potential of LLMs, with most past research focusing on improving single-step prompting. In the NLP field, structured languages like SQL pose a greater challenge than plain text. We introduce the workflow prompting paradigm to enhance the performance of LLMs in text-to-SQL tasks. By decomposing steps, reducing irrelevant information, and transforming the complex problem into a classification task, we can focus the attention of LLMs. Self-correction and active learning modules are designed according to the capability thresholds of LLMs, with different sub-modules performing their respective roles, further improving the upper limit of LLM-based prompting methods in text-to-SQL tasks. Adequate experiments on Spider, Spider-Realistic, and Bird datasets demonstrate that our approach significantly improves prompting performance, outperforming state-of-the-art fine-tuning and LLM-based methods. This illustrates the effectiveness of the workflow prompting paradigm in enhancing the problem-solving scope of LLMs.
Limitations
@@ Appendix B Prompts
B.0.1 Identify Problem Elements Prompt of Spider Dataset
@@
B.0.2 Identify Problem Elements Prompt of Bird Dataset
### Here are some examples:
query: Which year has the least number of movies that was released and what is the title of the movie in that year that has the highest number of rating score of 1?
hint: least number of movies refers to MIN(movie_release_year); highest rating score refers to MAX(SUM(movie_id) where rating_score = ’1’)
output: "query_entities": ["year", "number of movies that was released", "the title of the movie", "rating score"], "hint_entities": ["movie_release_year", "movie_id", "rating_score"], "query": "The year with the least number of movies released is identified, and the title of the movie with the highest rating score of 1 in that year is provided."
query: What’s the description of user 85981819’s movie list with the most followers?
hint: user 85981819 refers to user_id = 85981819; most followers refers to Max(list_followers); description refers to list_descriptions;
output: "query_entities": ["the description of user", "movie list", "followers"], "hint_entities": ["user_id", "list_followers", "list_descriptions"], "query": "The description of user 85981819’s movie list with the most followers is requested."
query: Who is the youngest player to have won the Purple Cap?
hint: Who refers to Player_Name; youngest player to have won the Purple Cap refers to min(subtract(Season_Year, DOB))
output: "query_entities": ["player", "the Purple Cap"], "hint_entities": ["Player_Name", "Season_Year", "DOB"], "query": "The name of youngest player to have won the Purple Cap."
How old is Ishan Kishan in 2022?
hint: old refers to SUBTRACT(2022, SUBSTR(DOB, 1, 4)); Ishan Kishan refers to Player_Name = ’Ishan Kishan’;
output: "query_entities": ["age", "Ishan Kishan", "2022"], "hint_entities": ["DOB", "Player_Name"], "query": "The age of Ishan Kishan in 2022"
query: What’s the average rating score of the movie "A Way of Life"?
hint: A Way of Life’ is movie_title; average rating score = Divide (Sum(rating_score), Count(rating_id));
output: "query_entities": ["rating score", "movie", "A Way of Life"], "hint_entities": ["movie_title", "rating_score", "rating_id"], "query": "The average rating score of the movie Ä Way of Life"̈
### Please be sure to follow the following specifications:
1."query_entities" refers to all entities in the query, including all description information in the query.
2."hint_entities" refers to all entities in the hint, including all description information in the hint.
3.Your output must be output in json format, and only this json needs to be returned. It needs to include all fields in json. The json format is as follows:
"query_entities":[query entities], "hint_entities":[hint entities], "query":"Rewritten question, removing unnecessary content"
query: [query input]
hint: [hint]
output:
B.1 Information Filter Prompt
@@ 3.2. The where statement condition only gives the conditions of the corresponding table;
B.2 Question Classification Prompt
@@
B.3 SQL Generation Prompt
B.3.1 Few-shot: easy class
@@
B.3.2 Few-shot: join class
@@
B.3.3 Few-shot: join-nested or nested class
@@
B.3.4 The Specification of Bird Dataset
1."LIMIT" is only used when explicitly requesting how much to retrieve in the query or for extreme value problems. 2.In sql, just select columns that are explicitly requested in the query. 3.Please note that the required fields in the ORDER BY clause and SELECT require the addition of a WHERE condition IS NOT NULL. 4.Note that a single-word field name does not require double quotes, but a field name composed of multiple words requires double quotes. 5.Pay attention to the division operation. It is necessary to perform a type conversion operation on molecule A, such as: cast(A as REAL).
B.4 Self-Correction Prompt
@@
B.5 Active Learning Prompt
B.5.1 Activate Learning Prompt of Spider Dataset
@@
B.5.2 Activate Learning Prompt of Bird Dataset
You are an experienced SQL engineer. Please correct the given problem according to the tips and examples given. If it is not within the tips range or there is no issues, return SQLite QUERY as is.
Tip1: In sql, just select columns that are explicitly requested in the query.
Question: Return the id and first name of the customer who has the fewest accounts.
SQLite SQL QUERY: SELECT T1.customer_id, T2.customer_first_name, T2.customer_last_name FROM Customers_cards AS T1 JOIN Customers AS T2 ON T1.customer_id = T2.customer_id GROUP BY T1.customer_id ORDER BY count(*) ASC LIMIT 1
Fixed SQL QUERY: SELECT T1.customer_id, T2.customer_first_name FROM Customers_cards AS T1 JOIN Customers AS T2 ON T1.customer_id = T2.customer_id GROUP BY T1.customer_id ORDER BY count(*) ASC LIMIT 1
Tip2: Pay attention to the columns that are used for the SELECT clause. Fix possible ambiguous columns if there are the same columns in different table in the SELECT clause and Do not use ‘||‘ to connect multiple fields.
Question: Please list the full name of the youngest person
SQLite SQL QUERY: SELECT first_name || ’ ’ || last_name AS full_name FROM member join profile on profile.id = member.id ORDER BY profile.age ASC LIMIT 1
Fixed SQL QUERY: SELECT member.first_name, member.last_name FROM member join profile on profile.id = member.id ORDER BY profile.age ASC LIMIT 1
Tip3: Pay attention to standard writing for optimal value problems.
Question: State the most popular movie?
HINT: most popular movie refers to MAX(movie_popularity);
SQLite SQL QUERY: SELECT movie_title FROM movies WHERE movie_popularity = (SELECT MAX(movie_popularity) FROM movies);
Fixed SQL QUERY: SELECT movie_title FROM movies ORDER BY movie_popularity DESC LIMIT 1
Tip4: Note that the joint table condition must be in the Foreign_key.
Foreign_key: [lists.user_id = lists_users.user_id, ratings.movie_id = movies.id, link.id = movies.id]; Question: What is the average rating for movie titled ’When Will I Be Loved’?
SQLite SQL QUERY: SELECT avg(ratings.rating_score) FROM movies INNER JOIN ratings ON movies.movie_id = ratings.movie_id WHERE movies.movie_title = ’When Will I Be Loved’
Fixed SQL QUERY: SELECT avg(ratings.rating_score) FROM movies INNER JOIN ratings ON movies.id = ratings.movie_id WHERE movies.movie_title = ’When Will I Be Loved’
The following is the question that requires you to decide whether to make changes
Database scheme: [table_info]
Question: [query][hint]
SQLite SQL QUERY: [sql]
Fixed SQL QUERY: SELECT